Find the roots of the polynomial equation. -x3 + 5x2 - 11x + 55 = 0
Answer by stanbon(75887)

Find the roots of the polynomial equation.
-x3 + 5x2 - 11x + 55 = 0
-----
I graphed it; x = 5 is a root.
-----------
Use Synthetic division to find the 2 other roots:
5)....-1....5....-11....55
......-1...0.....-11...|..0
-----
Quotient: -x^2 - 11 = 0
---
x = +sqrt(11), or x = -sqrt(11), or x = 5
=======================================================
Cheers,
Stan H.
===========